Price $11.00 $7.50 $6.00 $4.00H 0 4 6 9 12 15 QuantityDirections: Answer the questions below based on the graph.

  1. How many units will the domestic firms produce without trade?
  2. How many units will the domestic firms produce without a tariff if the foreign producer can sell the product at a $4 price?
  3. How many units will the foreign firms produce / sell if a government tariff of $2.00 is imposed on foreign goods?
  4. What will be the total government revenues if a tariff of $2.00 is imposed on foreign goods?
  5. What will be the total deadweight losses if a tariff of $2.00 is imposed on foreign goods?
Answer #1

a) Domestic firms will produce 9 units without trade

b) Domestic firms will produce 4 units when P=4

c) With a tariff, the price increases to 6 at which the firms will produce 6 units

d) Total government revenue = 2*imports where imports = QD-QS = 12-6 = 6

Revenue = 2*6 =12

e) DWL = 0.5*2*2 + 0.5*3*2

= 2+3

= 5